OpenStudy (anonymous):

A a population shrinks from its initial level of 24,000 at a continuous decay rate of 5.9% per year. (a) Find a formula for P(t), the population in t years. (b) By what percent does the population shrink each year? I found the formula: P=24,000e^(-0.059t) for (a) but I have no idea how to find (b)

OpenStudy (anonymous):

@ganeshie8 you can explain here if its easier

ganeshie8 (ganeshie8):

^^interpret that

OpenStudy (anonymous):

I get where you get the e^0.059 but I don't understand why you multiplied it by -1

ganeshie8 (ganeshie8):

Okay, try this : Population in the start = 24,000 Population after one year = 24,000e^(-0.059(1)) = 24,000e^(-0.059)

ganeshie8 (ganeshie8):

evaluate them both, and find the difference

ganeshie8 (ganeshie8):

that tells you how many ppl (population) decayed/dead in first year

ganeshie8 (ganeshie8):

difference = 24,000e^(-0.059) - 24,000 = 24,000(e^(-0.059) -1 )

ganeshie8 (ganeshie8):

divide it by the start amount to get the percent change

ganeshie8 (ganeshie8):

24,000(e^(-0.059) -1 ) ------------------ = e^(-0.059) - 1 24000

OpenStudy (anonymous):

oh okay that makes more sense. In wolfram it looked like you were doing e^(-.0.059*-1) and I didn't understand

ganeshie8 (ganeshie8):

ahh wolfram is somhow making it in unreadable format lol http://www.wolframalpha.com/input/?i=%28e%5E%28-0.059%29%29++-+++1

OpenStudy (anonymous):

lol yeah wolfram can be touchy. but it makes sense just fine without it. I do get it now

ganeshie8 (ganeshie8):

good :)